# GlucoLens: Explainable Postprandial Blood Glucose Prediction from Diet and Physical Activity

### v1

@@ Abstract
Postprandial hyperglycemia, marked by the blood glucose level exceeding the normal range after meals, is a critical indicator of progression toward type 2 diabetes in prediabetic and healthy individuals. A key metric for understanding blood glucose dynamics after eating is the postprandial area under the curve (PAUC). Predicting PAUC in advance based on a person’s diet and activity level and explaining what affects postprandial blood glucose could allow an individual to adjust their lifestyle accordingly to maintain normal glucose levels. In this paper, we propose GlucoLens, an explainable machine learning approach to predict PAUC and hyperglycemia from diet, activity, and recent glucose patterns. We conducted a five-week user study with 10 full-time working individuals to develop and evaluate the computational model. Our machine learning model takes multimodal data including fasting glucose, recent glucose, recent activity, and macronutrient amounts, and provides an interpretable prediction of the postprandial glucose pattern. Our extensive analyses of the collected data revealed that the trained model achieves a normalized root mean squared error (NRMSE) of 0.123. On average, GlucoLense with a Random Forest backbone provides a 16% better result than the baseline models. Additionally, GlucoLens predicts hyperglycemia with an accuracy of 74% and recommends different options to help avoid hyperglycemia through diverse counterfactual explanations. Code available: https://github.com/ab9mamun/GlucoLens.
Index Terms:
@@
machine learning, metabolic health, continuous glucose monitoring, diabetes, hyperglycemia
I Introduction
@@
Hyperglycemia, or high blood sugar, occurs when the body cannot effectively regulate glucose levels. Hyperglycemia is defined as having a blood glucose (BG) level above 140 mg/dL [1]. Lack of physical activity and overconsumption of carbohydrates can affect a person’s metabolism and affect the ability to regulate glucose [2]. People with untreated hyperglycemia are at increased risk of developing complications such as eye problems (retinopathy), kidney problems (nephropathy), nerve damage (neuropathy), heart disease, stroke, poor blood flow to the limbs, and are also more likely to experience depression [3]. While hyperglycemia can affect anyone, individuals with prediabetes or diabetes are at a higher risk compared to healthy people [4, 5].
The Centers for Disease Control (CDC) estimates that 38% of American adults are prediabetic and almost 20% of them are unaware of their condition [6]. To help prevent the increasing prevalence of hyperglycemia and prediabetes, the Food and Drug Administration (FDA) has approved the sale of Continuous Glucose Monitors (CGMs) over the counter in the U.S. in 2024, making them more accessible to people with or without diabetes [7]. Prediabetes can be reversed with proper lifestyle management, such as diet and physical activity [8]. However, if left untreated, it can develop type 2 diabetes, which is an irreversible lifelong condition [9].
@@
Figure 1: An example of calculation of Postprandial Area Under the Curve (PAUC), Increase in PAUC from the projected area after lunch ($\Delta$PAUC), and Maximum Postprandial Blood Glucose (MaxPBG).
@@
Healthy individuals are expected to maintain their BG between 60 mg/dL and 140 mg/dL, but it is not uncommon for them to have BG over the recommended range for a short duration after eating certain meals. The area under the BG curve for a certain duration (e.g., 2 or 3 hours) after a meal is known as the Postprandial Area Under the Curve (PAUC). An example of calculating PAUC, change in PAUC, and the maximum postprandial blood glucose (MaxPBG) is shown in Fig. 1. The value of the PAUC is an important indicator of healthy BG regulation. Being obese or overweight, or in general, a higher body mass index (BMI) has been linked to having a higher risk of developing diabetes [10] in the previous studies. Blood glucose forecasting is an active area of research that has been addressed with different machine learning and deep learning approaches including ensemble methods, attention methods, and knowledge distillation [11, 12, 13]. However, to the best of our knowledge, the use of diet, physical activity, work routine, and previous baseline glucose parameters to predict the area under the glucose curve remains unexplored. Despite the increasing interest in the area under the glucose curve as a metric to estimate the risk of hyperglycemia [14, 15], an in-depth ML-based analysis is yet to be published.
In this paper, we aim to address quantified relationships of diet, physical activity, and work routine with hyperglycemia. Moreover, we propose a machine learning solution for estimating the postprandial AUC based on easily available information such as CGM readings, diet, work routine, and wearable sensor data. To summarize, our contributions are as follows: (i) formulating a problem of predicting PAUC in advance based on diet, work habit, and physical activity, (ii) designing a system with wearables and log-keeping to solve the problem, (iii) conducting a user study to collect the required data, and (iv) implementing GlucoLens for PAUC and hyperglycemia prediction and diverse counterfactual (CF) explanations.
II Materials and Methods
II-A System Overview
To investigate the questions about the relationships of diet and activity with postprandial BG, we designed a system of four data sources: a wearable device with motion sensors, a CGM device, food logs, and work logs. As this study was not only about prediction but also about interpretability, the system’s data processing tool had to be able to process the raw data and convert them to compatible electronic formats, which would also be human-readable. The processed data would then be used to train ML models which are capable of making accurate predictions about the PAUC given the type of diet and activity. The system is illustrated in Fig. 2.
@@ II-A System Overview
Figure 2: The GlucoLense system’s body sensor network is composed of food logs, work logs, a CGM device, and a wearable sensor. Data from all these sources are combined to create a unified dataset to train a model.
Suppose, for a specific day $i$, we have food logs $X_{food}^{(i)}$, work logs $X_{work}^{(i)}$, wearable data $X_{wearabele}^{(i)}$, and CGM readings $X_{CGM}^{(i)}$, then our goal is to learn the function $f_{y}$ so that it reads the logs and provides the estimated outcome, $\hat{y}$, which is very close to the actual outcome, $y$.
$\hat{y}=f_{y}(X_{food}^{(i)},X_{work}^{(i)},X_{CGM}^{(i)},X_{wearabele}^{(i)})$
II-B User Study
@@
To develop the proposed system, we conducted a user study by recruiting full-time working individuals living in a U.S. city. Appropriate IRB approval and informed consent were obtained before the data collection. Each participant was given a CGM device and an activPAL wearable [16] device. Lunch was delivered at work from select restaurants on every working day and each participant received printed forms to maintain the food logs and work logs. The option of homemade meals was not considered for lunch on working days because it would make the task of estimating the macronutrients difficult. On the other hand, the selected restaurants already had the macronutrients for all their items available online, so it was comparatively easier to estimate the nutritional components for every meal consumed.
@@
In the food logs, the participants noted down the time and what they ate in every meal including if there were any leftovers. The participants also logged the amount of water and any supplements (e.g., multivitamins) they were taking every day. In the work logs, they noted when they started working, when they stopped working, how they arrived at work or if they were working from home, and approximately what percentage of working time they spent sitting, standing, and walking.
@@
The data collection ran in three different periods: a 1-week baseline period, a 2-week intervention period named ‘Condition 1’, and a 2-week intervention period named ‘Condition 2’. The baseline week did not have any intervention but the participants wore their CGM devices and activPAL. Two interventions were ‘recommending to be standing’ and ‘recommending to be moving’ as much as possible during work. All participants were provided with both of the interventions, but the temporal order was random. Therefore, for some participants, the temporal order was the baseline, stand, move, and for the other participants, it was the baseline, move, stand. In this study, we have included our analysis based on the data of 10 participants having an average baseline BMI of $32.8\pm 4.5$.
The activPAL sensor records an event whenever the user’s activity changes. The different activities or events detected by the device are: sedentary, standing, stepping, cycling, primary lying, secondary lying, and seated transport. For each event, the event start time and the duration of the event are recorded. The CGM device reported the blood glucose level averaged over every 15-minute window.
II-C Data Processing
@@
In this study, one of the main interests was predicting postprandial blood glucose during work days and investigating how diet and activity during work days affect postprandial AUC. Moreover, the participants were provided standardized lunch on the working days, whereas they ate anything they liked on the weekends. In the food logs, the participants kept track of any leftover portions which allowed us to accurately extract the amount of calories and macronutrients consumed during lunch. The food logs were maintained as handwritten logs. They were processed through Google Cloud Vision OCR to create electronic logs followed by some human intervention for issues that could not be resolved by the OCR. The amounts of macronutrients consumed were used to estimate glycemic loads (GLs) using the formula from [17].
@@
The work logs were also handwritten and they were processed manually to convert them to electronic format. From activPAL sensors data and with the help of work logs, the durations of sitting, standing, and stepping were calculated for the day to lunch, as well as the durations of sitting, standing, and stepping during work hours for the day until lunch. For this work, fasting glucose was defined as the minimum CGM reading between 6 AM and 10 AM. The recent CGM was defined as the average CGM reading of the same prediction day from midnight to 8 AM. A complete list of the input features of the ‘Sensor+Macro’ feature set can be found in Fig. 3. Two additional inputs are used in the ‘All’ feature set: an activity score calculated from self-reported activity logs during work and the glycemic load. In the ‘Self+Macro’ feature set, the 6 features containing sitting, standing, and stepping durations are replaced by the activity score based on self-reported activity logs. The activity score was calculated from the self-reported activities in the work logs. In the work logs, each user reported the percentage of their working hours spent sitting, standing, and walking. As mentioned in Section II-B, there were three phases: Baseline, Condition 1, and Condition 2. The recent activity score is calculated by taking the average percentage of time spent in walking activity in the previous days of the same phase and adding with $\frac{1}{2}\times$ the average percentage of time spent in standing activity in the previous days of the same phase. In the ‘Self+GL’ and ‘Sensor+GL’ feature sets, the macronutrients: net carbs, fat, protein, and fiber are removed and replaced with the glycemic load calculated in Equation 1. The relationship of all five different feature sets can be found in Fig. 3.
Figure 3: A comparison of the five different feature sets used in this study. GL = glycemic load of the meal. Self = activity scores calculated by self-reported activity duration in work logs. Macro = macronutrients that are used to calculate glycemic load: net carb, fat, protein, and carb. Sensor = activity metrics from activPAL sensor: duration of sitting, standing, and stepping activities of the day before lunch and the same metrics from the start of working to just before lunch of the same day.
TABLE I: Size of the dataset depending on the feature set chosen. The meanings of the different feature sets are explained in Fig. 3.
Feature Sensor Self Sensor Self All
set +Macro +Macro +GL +GL
Train size 127 140 127 140 127
Test size 32 35 32 35 32
II-D ML Models
Three regression models, Random Forest (RF), Ridge Regressor, and MLP Regressor were trained for each of the three tasks: estimating PAUC, $\Delta$PAUC, and MaxPBG. MLPs and Random Forests are capable of making effective predictions with proper feature extractions [18, 19], whereas linear regression methods with regularization, such as Ridge Regression [20] and Lasso Regression [21] are popular choices as baselines. Neural networks are usually data-hungry, which is why, on small datasets, classical ML methods can show competitive and even better performance than neural networks.
Three variations of RF, three variations of Ridge, and thirteen variations of MLP were trained and tested. Therefore, in total 19 models, 3 tasks each, 5 feature sets each, and 3 random seeds were used. The variations of RF, Ridge, and MLP can be found in Table II and Table LABEL:tbl:mlp_hyperparams. For each experiment, 20% data instances were assigned to the test set, and 80% data instances were assigned to the training set. The exact number of instances for training and test sets are represented in Table I.
II-E Hyperglycemia Detection and Counterfactual Explanations
GlucoLens was also trained for hyperglycemia detection with RF backbone models. Both MLP and RF models were trained for hyperglycemia detection and RF (avg accuracy 74%) outperformed MLP (best accuracy 71%). Counterfactual explanations can provide insights on features responsible for an undesired health outcome and possible remedies to overcome them [22, 23]. As a primary goal of the paper was to explore more knowledge about the reasons for and ways to prevent hyperglycemia, a DiCE-based counterfactual explanations generator is integrated with GlucoLens. GlucoLens provides multiple counterfactual explanations that are diverse and achievable with small perturbations from the original example [24].
TABLE II: An overview of all experiments conducted for the regression problem. MLP = multilayer perception. RF = random forest. Ridge = ridge regression. $n_{est}$ = number of estimators in random forest. Meanings of the outcomes PAUC, $\Delta$PAUC, and MaxPBG can be found in Fig. 1.
Target outcomes PAUC, $\Delta$PAUC, MaxPBG

### v2

@@ Abstract
Postprandial hyperglycemia, marked by the blood glucose level exceeding the normal range after consuming a meal, is a critical indicator of progression toward type 2 diabetes in people with prediabetes and in healthy individuals. A key metric for understanding blood glucose dynamics after eating is the postprandial area under the curve (AUC). Predicting postprandial AUC in advance based on a person’s lifestyle factors, such as diet and physical activity level, and explaining the factors that affect postprandial blood glucose could allow an individual to adjust their lifestyle accordingly to maintain normal glucose levels. In this study, we developed an explainable machine learning solution, GlucoLens, that takes sensor-driven inputs and uses advanced data processing, large language models, and trainable machine learning models to predict postprandial AUC and hyperglycemia from diet, physical activity, and recent glucose patterns. We used data obtained from wearables in a five-week clinical trial of 10 adults who worked full-time to develop and evaluate the proposed computational model that integrates wearable sensing, multimodal data, and machine learning. Our machine learning model takes multimodal data from wearable activity and glucose monitoring sensors, along with food and work logs, and provides an interpretable prediction of the postprandial glucose pattern. Our GlucoLens system achieves a normalized root mean squared error (NRMSE) of 0.123 in its best configuration. On average, the proposed technology provides a 16% better performance level compared to the comparison models. Additionally, our technique predicts hyperglycemia with an accuracy of 73.3% and an F1 score of 0.716 and recommends different treatment options to help avoid hyperglycemia through diverse counterfactual explanations. Code available: https://github.com/ab9mamun/GlucoLens.
keywords:
@@
machine learning, metabolic health, continuous glucose monitoring, diabetes, hyperglycemia, large language models
1 Introduction
Wearable sensors have become indispensable tools in the field of healthcare and personalized medicine, offering continuous, non-invasive monitoring of physiological parameters in real time Vaghasiya et al. (2023); Clubbs Coldron et al. (2023). Devices such as continuous glucose monitors (CGMs), physical activity trackers, and smartphone sensors provide valuable insights into an individual’s metabolic health, physical activity, and lifestyle habits Uddin and Soylu (2021); Mamun et al. (2022). These sensors enable the collection of granular data that can be used to understand complex physiological processes, predict health outcomes, and design interventions Shields et al. (2024); Steck et al. (2019). Wearable sensors also provide opportunities for developing robust solutions despite different challenges, such as small dataset, missing features, data in the wild, etc. Mamun et al. (2022); Maray et al. (2023); Mamun et al. (2025); Soumma et al. (2025). In the context of diabetes management, wearable sensors allow for the precise tracking of blood glucose fluctuations and physical activity levels, offering individuals and clinicians the means to implement timely and tailored strategies for disease prevention and management Pickup et al. (2015). This accessibility to continuous health data marks a significant shift toward proactive and data-driven healthcare solutions.
@@ 1 Introduction
Hyperglycemia, or high glucose concentration in the blood, occurs when the body cannot effectively regulate glucose levels. Hyperglycemia in the fasted state (i.e., no caloric intake for at least 8 hours) is defined by a fasting plasma glucose level $\geq$ 100 mg/dL ($\geq$5.6 mmol/L), whereas hyperglycemia in the non-fasted state generally is defined as having a blood glucose level (BGL) $\geq$140 mg/dL (7.8 mmol/L) two hours after a meal nih ; Gehlaut et al. (2015). Lack of physical activity and relative overconsumption of carbohydrates can affect a person’s metabolism and their ability to regulate glucose Crichton and Alkerwi (2015). Other common risk factors often cited as potentially responsible for developing diabetes are being obese, overweight, or having a higher than normal body mass index (BMI) Maggio and Pi-Sunyer (2003). Untreated hyperglycemia increases the risk of complications like retinopathy, nephropathy, neuropathy, cardiovascular disease, stroke, poor limb circulation, and depression Mouri and Badireddy (2023). While hyperglycemia can affect anyone, individuals with prediabetes or diabetes are at a higher risk compared to healthy people Alvarez et al. (2017); Castro (2022).
@@
Figure 1: An example of postprandial Area Under the Curve (AUC), incremental AUC (iAUC), and Maximum postprandial Blood Glucose Level (MaxBGL) for one participant after a lunch meal.
The Centers for Disease Control (CDC) estimates that 38% of American adults are prediabetic and 19% of them are unaware of their condition CDC . To help prevent the increasing prevalence of hyperglycemia and prediabetes, the Food and Drug Administration (FDA) has approved the sale of Continuous Glucose Monitors (CGMs) over the counter in the U.S. in 2024 fda . This decision has made CGM devices more accessible to people with or without diabetes. Prediabetes can be reversed with proper lifestyle management, such as diet and physical activity Echouffo-Tcheugui et al. (2023). However, if left untreated, it can develop type 2 diabetes, which is an irreversible lifelong condition Chakarova et al. (2019). Healthy individuals are generally expected to maintain their blood glucose levels (BGL) within a range of 60 to 140 mg/dL nih . However, temporary elevations above this range are common following certain meals. The area under the BGL curve during a specific post-meal period (e.g., 2–3 hours) is referred to as the postprandial Area Under the Curve (AUC), which is an important indicator of healthy BGL regulation. Numerous studies have explored methods for calculating AUC and the various factors influencing it Floch et al. (1990); Sakaguchi et al. (2016); T GALE et al. (2023). An example of calculating AUC, incremental AUC, and the maximum postprandial blood glucose level (MaxBGL) is shown in Fig. 1.
@@
Blood glucose forecasting using wearable sensor data is an active area of research that has been addressed with different machine learning and deep learning approaches, including ensemble methods, attention methods, and knowledge distillation Han et al. (2024); Farahmand et al. (2024, 2025). Despite the increasing interest in the area under the glucose curve as a metric to estimate the risk of hyperglycemia Uemura et al. (2017); Ugi et al. (2016), to the best of our knowledge, the use of diet, physical activity, work routine, and previous baseline glucose parameters to predict hyperglycemia and the area under the glucose curve remains unexplored.
In this paper, we leverage wearable sensor data and propose and develop a machine learning solution, GlucoLens, for estimating the postprandial AUC and predicting hyperglycemia with wearable sensor data from CGM and activity trackers, as well as food and work logs. To summarize, our contributions are as follows: (i) formulating a problem of predicting AUC and hyperglycemia in advance based on diet, work habit, and physical activity, (ii) designing hyperglycemia prediction systems based on machine learning and large language models that combines wearable data, food logs, work logs, and recent CGM patterns to make prediction, (iii) implementing GlucoLens for AUC and hyperglycemia prediction and diverse counterfactual (CF) explanations, (iv) train and test our models on a dataset collected in a clinical trial conducted by our team, (v) provide a thorough discussion of our findings and how it advances the research of hyperglycemia prediction.
2 Materials and Methods
((a))
((b))
@@ 2 Materials and Methods
Figure 2: The GlucoLens system’s body sensor network is composed of food logs, work logs, a CGM device, and a wearable sensor. Data from all these sources are combined to create a unified dataset to train a model. The machine learning models are further supported by large language models (LLM) and data augmentation tools based on Gaussian noise. Finally, counterfactual explanations are provided so that interventions can be applied.
2.1 System Overview
To investigate the questions about the relationships of diet and activity with postprandial BGL, we designed a system of four data sources: a wearable device with motion sensors, a CGM device, food logs, and work logs. The system’s data processing tools process the raw data from multiple sources and convert them to compatible electronic formats. The processed data are then used to train ML models which are capable of making accurate predictions about the AUC and hyperglycemia given the type of diet and activity. The system is illustrated in Fig. 2.
((a))
((b))
((c))
Figure 3: Wearables with sensors in the WorkWell Study. activPAL uses a tri-axial accelerometer for detecting activities. The FreeStyle Libre CGM patch measures glucose levels in the interstitial fluid (the fluid surrounding cells) using a thin filament inserted under the skin. The GENEActiv device uses a 3-axis accelerometer, a light sensor (photodiode), and a temperature sensor.
The system has three wearable devices: a FreeStyle Libre CGM device, an activPAL micro device, and a GENEActiv wearable wristband, as shown in Fig. 3. activPAL provides event data with the timestamps of that specific event pal . Changes in activity are considered as events by the activPAL device, e.g, when a person stands up from a sedentary position or starts moving from a standing position are considered events. The different activities or events detected by the device are: sedentary, standing, stepping, cycling, primary lying, secondary lying, and seated transport. For each event, the event start time and the duration of the event are recorded. Duration spent in different activities can be calculated from those events. The Freestyle Libre provides CGM readings in 15 minute-intervals. GENEActiv provides various biomarkers including acceleration, physical activity intensity, sedentary vs movement activity, sleep wake time act .
GlucoLens is composed of a multimodal data processing package, interfaces for an array of large language models (LLMs), trained machine learning models curated for postprandial AUC and hyperglycemia predictions, and a counterfactual explanation generator. The data processing package includes Optical Character Recognition (OCR), food log processor, CGM processor, activity processor, and work log processor. Our system processes the raw data from the sensors and these sources and prepare a multimodal dataset. The multimodal data is used to produce predictions from 7 different LLMs: GPT-3.5-Turbo, GPT-4, Claude Opus 4, Deepseek V3, Gemini Flash 2.0, Grok 3, and Mistral Large. The Mistral Large model declined to provide predictions, hence omitted from the results, and was not considered for the hybrid predictors. Additionally, machine learning backbones such as Random Forest (RF), Ridge Regressor, Multilayer Perceptron (MLP) Regressor, as well as XGBoost Chen and Guestrin (2016), and TabNet Arik and Pfister (2021). The multimodal dataset is augmented by adding small Gaussian noise to the scaled features. The system then produces predictions for postprandial AUC, maximum blood glucose level, and postprandial hyperglycemia. Finally, diverse counterfactual explanations are provided so that interventions can be designed to avoid hyperglycemia.
2.2 Wearables and Lifestyle Data
Figure 4: Timeline of the WorkWell Study. Each participant completed a 1-week Baseline phase data collection, 2 weeks of Condition1 phase data collection, and 2 weeks of Condition2 phase data collection. There was a 1-week break between Condition 1 and Condition 2. In all three phases, participants provided CGM, activPAL, GENEActiv, food logs, and work logs data. The Baseline phase was the control phase, where the participants maintained their usual work habits. Participants were randomly assigned an intervention of ‘Stand’ or ‘Move’ in Condition 1, and the other intervention was assigned in Condition 2. In the ‘Stand’ intervention, they were told to stand as much as possible at work, and during the ‘Move’ intervention, they were told to move as much as possible at work.
@@ 2.2 Wearables and Lifestyle Data
To develop the proposed system, we used a subset of data obtained in the WorkWell Study Wilson et al. (2023), a clinical trial (NCT04269070) conducted by our team that involved adults who worked full time. The WorkWell Study was approved by the Arizona State University Institutional Review Board, and all participants provided written informed consent to participate. Each participant was given a CGM device and an activPAL wearable Lyden et al. (2017) device. Lunch was delivered at work from select restaurants on every working day (five days per week) for one work week at Baseline, two work weeks during Condition 1, and two work weeks during Condition 2, as shown in Fig. 4. Each participant received printed forms to maintain the food logs and work logs. The option of homemade meals was not considered for lunch on working days because it would make the task of estimating the macronutrients difficult. The selected restaurants already had the macronutrients for all their items available online, which was used to estimate the nutritional components for every meal consumed.
@@
In the food logs, the participants noted down the time and what they ate in every meal, including if there were any leftovers. The participants also logged the amount of water and any supplements (e.g., multivitamins) they were taking every day. In the work logs, they noted when they started working, when they stopped working, how they arrived at work or if they were working from home, and approximately what percentage of working time they spent sitting, standing, and walking.
@@
The Baseline week did not have any intervention but the participants wore their CGM devices and activPAL. Two interventions for Condition 1 and Condition 2 were ‘recommending to be standing’ and ‘recommending to be moving’ as much as possible during work. All participants were provided with both interventions, but the temporal order was random. Therefore, for some participants, the temporal order was the Baseline, Stand, Move, and for the other participants, it was the Baseline, Move, Stand. In this study, we have included our analysis based on the data of 10 participants having an average Baseline BMI of $32.8\pm 4.5$.
2.3 Data Processing
Table 1: A comparison of the five different feature sets used in this study. The checkboxes indicate whether a feature is present in a specific feature set. GL = glycemic load of the meal. Self = activity scores calculated by self-reported activity duration in work logs. Macro = macronutrients that are used to calculate glycemic load: net carb, fat, protein, and carb. Sensor = activity metrics from the activPAL sensor: duration of sitting, standing, and stepping activities of the day before lunch and the same metrics from the start of working to just before lunch of the same day.
No. Feature name/ Sensor Sensor Self Self All
shorthand +GL +Macro +GL +Macro
1 Fasting glucose $\boxtimes$ $\boxtimes$ $\boxtimes$ $\boxtimes$ $\boxtimes$
2 Recent CGM $\boxtimes$ $\boxtimes$ $\boxtimes$ $\boxtimes$ $\boxtimes$
3 Lunch time $\boxtimes$ $\boxtimes$ $\boxtimes$ $\boxtimes$ $\boxtimes$
4 Work from home $\boxtimes$ $\boxtimes$ $\boxtimes$ $\boxtimes$ $\boxtimes$
5 BMI $\boxtimes$ $\boxtimes$ $\boxtimes$ $\boxtimes$ $\boxtimes$
6 Calories $\boxtimes$ $\boxtimes$ $\boxtimes$ $\boxtimes$ $\boxtimes$
7 Calories from fat $\boxtimes$ $\boxtimes$ $\boxtimes$ $\boxtimes$ $\boxtimes$
8 Saturated fat $\boxtimes$ $\boxtimes$ $\boxtimes$ $\boxtimes$ $\boxtimes$
9 Trans fat $\boxtimes$ $\boxtimes$ $\boxtimes$ $\boxtimes$ $\boxtimes$
10 Cholesterol $\boxtimes$ $\boxtimes$ $\boxtimes$ $\boxtimes$ $\boxtimes$
11 Sodium $\boxtimes$ $\boxtimes$ $\boxtimes$ $\boxtimes$ $\boxtimes$
12 Total carbs $\boxtimes$ $\boxtimes$ $\boxtimes$ $\boxtimes$ $\boxtimes$
13 Sugar $\boxtimes$ $\boxtimes$ $\boxtimes$ $\boxtimes$ $\boxtimes$
14 Work start time $\boxtimes$ $\boxtimes$ $\boxtimes$ $\boxtimes$ $\boxtimes$
15 Day of the week $\boxtimes$ $\boxtimes$ $\boxtimes$ $\boxtimes$ $\boxtimes$
16 activPAL $\boxtimes$ $\boxtimes$ $\square$ $\square$ $\boxtimes$
17 Self reported acitivity $\square$ $\square$ $\boxtimes$ $\boxtimes$ $\boxtimes$
18 GL $\boxtimes$ $\square$ $\boxtimes$ $\square$ $\boxtimes$
19 Net carbs $\square$ $\boxtimes$ $\square$ $\boxtimes$ $\boxtimes$
20 Fat $\square$ $\boxtimes$ $\square$ $\boxtimes$ $\boxtimes$
21 Protein $\square$ $\boxtimes$ $\square$ $\boxtimes$ $\boxtimes$
22 Fiber $\square$ $\boxtimes$ $\square$ $\boxtimes$ $\boxtimes$
@@
In this study, one of the main interests was predicting postprandial blood glucose during work days and investigating how diet and activity during work days affect postprandial AUC. Moreover, the participants were provided standardized lunch on working days, whereas they ate anything they liked on weekends. In the food logs, the participants kept track of any leftover portions, which allowed us to accurately extract the amount of calories and macronutrients consumed during lunch. The food logs were maintained as handwritten logs. They were processed through Google Cloud Vision OCR to create electronic logs, followed by some human intervention for issues that could not be resolved by the OCR. The amounts of macronutrients consumed were used to estimate glycemic loads (GLs) using the formula from Lee et al. (2021) as shown in Equation 1.
@@
The work logs were also handwritten, and they were processed manually to convert them to an electronic format. From activPAL sensors data and with the help of work logs, the durations of sitting, standing, and stepping were calculated for the day to lunch, as well as the durations of sitting, standing, and stepping during work hours for the day until lunch. For this work, fasting glucose was defined as the minimum CGM reading between 6 AM and 10 AM. The recent CGM was defined as the average CGM reading of the same day from midnight to 8 AM. A complete list of the input features of the different feature sets that were formed can be found in Table 1. The ‘Sensor+Macro’ feature set uses activity metrics from the sensor data and calculated macronutrients, along with the common features that are present in all feature sets. Two additional inputs are used in the ‘All’ feature set: an activity score calculated from self-reported activity logs during work and the glycemic load. In the ‘Self+Macro’ feature set, the 6 features containing sitting, standing, and stepping durations are replaced by the activity score based on self-reported activity logs. The activity score was calculated from the self-reported activities in the work logs. In the work logs, each participant reported the percentage of their working hours spent sitting, standing, and walking. The recent activity score is calculated by taking the average percentage of time spent in walking activity in the previous days of the same phase and adding with $\frac{1}{2}\times$ the average percentage of time spent in standing activity in the previous days of the same phase. In the ‘Self+GL’ and ‘Sensor+GL’ feature sets, the macronutrients: net carbs, fat, protein, and fiber are removed and replaced with the glycemic load calculated in Equation 1. The relationship of all five different feature sets can be found in Table 1.
2.4 ML Models
Table 2: An overview of the different backbones and their hyperparameters for the regression problem. Explanations of the feature sets can be found in Table 1. MLP = multilayer perceptron. RF = random forest. Ridge = ridge regression. $n_{est}$ = number of estimators in random forest. Meanings of the outcomes AUC and MaxBGL can be found in Fig. 1. Gly_Hybrid = best performing regressor (RF) with support of LLM-based predictions, Gly_Hybrid_v2 is RF with LLM-based predictions but with only the best LLM (Claude Opus 4). Gly_Max is Gly_Hybrid_v2 with augmented training data with Gaussian noise.
Target outcomes AUC, MaxBGL, Hyperglycemia